A random sample of 20 family practitioners in a city had an average salary of $150,250 with a standard deviation of $32,400. Can you conclude that the true average family practitioner salary in this city is less than $180,000? It has been confirmed that the these salaries come from a normal distribution.

What is the test statistic? Round to the fourth as needed.
Answer #1

Solution:

The null and alternative hypothesis is ,

H0 :  \mu = $180,000

Ha : \mu < $180,000

Test statistic (t) =

= (\bar x - \mu ) / s / \sqrt n

= (150,250 - 180,000) / 32,400 / \sqrt 20

Test statistic = -4.11